# JWST Insights into Narrow-line Little Red Dots

### v1

@@ 3.2 Selection of Narrow-line LRDs
A broad Gaussian component (FWHM 1000 ) and a narrow Gaussian component (FWHM 1000 ).
@@
The [N ii] lines are added if needed, and they are forced to have the same width as the narrow component of H. To compare the quality of these two types of fitting, we calculate the Bayesian Information Criterion (BIC) parameter, defined as (A. R. Liddle, 2007):
@@
Based on the above procedure, we identify five narrow-line LRDs with no broad H line component from the 32 sources. We also attempt to fit the H lines of these five sources using two narrow Gaussian components (FWHM 1000 ). The resultant BIC value still favors the single narrow component model in each case. The fraction of narrow-line LRDs in our sample is . This fraction may be biased by the complex selection process of the NIRSpec observations. On the other hand, we argue that the photometric selection is not sensitive to the line width, and thus the narrow-line LRD fraction is unlikely affected significantly.
@@ 3.3 Morphology
The NIRCam images of the five narrow-line LRDs in Figure 4 show that they are very compact. To quantitatively measure their morphology and detect their potential extended emission, we fit their multi-band JWST images (Figure 5) using GalfitM (B. Häußler et al., 2013; M. Vika et al., 2013), a multi-band version of the two-dimensional image fitting code Galfit (C. Y. Peng et al., 2002, 2010). We first cut images for each band from the background-subtracted mosaic images. The corresponding error cutout images are used as input sigma images. To ensure that the best-fit value is close to unity, we scale each sigma image by a factor of – such that its median background pixel value matches the standard deviation of the background pixels in the science image. Bands with no significant source detection are excluded during the fitting. For UID-16321, F090W and F115W are excluded; for UID-20547, F090W, F115W, and F150W are excluded; and for UID-8219 and UID-22015, F090W is excluded.
@@
It is still unclear whether LRDs are AGN-dominated or galaxy-dominated, particularly in the case of narrow-line LRDs. Therefore, we perform three types of model fitting: (1) a single PSF model, representing an AGN-dominated morphology; (2) a single Sérsic model, representing a pure galaxy; and (3) a combination of a PSF and a Sérsic component, representing an AGN and its host galaxy. The PSF models used are those constructed in Section 2.1. We fit a region centered on the source. During the fitting procedure, the axis ratio, half-light radius, and Sérsic index of the galaxy component are allowed to vary following a -order Chebyshev polynomial, while the position angle is allowed to vary following a -order Chebyshev polynomial. These flexibilities serve to accommodate the morphological difference between long and short wavelengths, as these sources are often more extended in shorter wavelengths.
@@
For UID-16321, UID-20547, and UID-22015, their multi-band JWST image can be well fitted by a single PSF model. UID-16321 and UID-20547 are very faint in the short wavelengths, so only their long-wavelength data (where LRDs typically appear unresolved) are used to measure the morphology. For UID-20547, it F277W band residual of the PSF fitting suggests weak extended emission. For UID-8219 and UID-12329, their short-wavelength images cannot be well fitted by a single PSF model, revealing clear extended emission. UID-8219 can be sufficiently fitted with a Sérsic component, and an additional PSF component slightly decreases the value from 1.13 to 1.09. In case of UID-12329, a single Sérsic model also fails to provide a good fit, with significant residuals in the short-wavelength bands. We thus adopt a PSF+Sérsic model that yields a reasonable result with value of 1.24. Figure 5 shows the PSF+Sérsic fitting result for UID-12329. The fraction of the Sérsic component decreases toward longer wavelengths. The Sérsic component of UID-12329 is slightly offset from the central point sources; such offset extended emission are also seen in other LRDs (e.g., C.-H. Chen et al., 2025b, a).
3.4 Emission-line Properties
@@
Using the emission line quantities measured in Section 3.2, we analyze the emission line properties of the narrow-line LRDs and compare them with those of normal star-forming galaxies. We adopt galaxy samples from K. Nakajima et al. (2023) and M. Curti et al. (2024) and measure their properties using the same method as applied to our sources. Figure 6 presents a comparison of the observed FWHM, intrinsic FWHM, luminosity, and equivalent width (EW) of the [O iii], H, and H lines among narrow-line LRDs, broad-line LRDs, and the star-forming galaxies. For broad-line LRDs, we only use the narrow components of the H and H lines for comparison. UID-16321 does not have medium- or high-resolution spectral coverage of the [O iii] line, thus its values are not available. We measure its from the prism spectrum. The comparison of the FWHM between the [O iii] and H lines in Figure 6(d) does not account for the LSF effect, as these two lines have similar wavelengths and spectral resolutions.
@@
Figure 6 suggests that narrow-line LRDs may exhibit different emission-line properties compared to typical star-forming galaxies. For star-forming galaxies, their observed are slightly smaller than (Figure 6a), because the NIRSpec grating spectrum has higher spectral resolutions at longer wavelengths. After correcting the LSF effect, is consistent with (Figure 6b). For the narrow-line LRDs, three of them have slightly larger than . On the other hand, three LRDs have medium-resolution spectra covering H and they all show consistent and . UID-22015 shows tentative larger but its measurement error is . The observed of narrow-line LRDs also tend to be larger compared with the star-forming galaxies with similar as shown in Figure 6(c). Conversely, the narrow-line LRDs generally exhibit smaller H EWs and larger [O iii] EWs compared with star-forming galaxies, although two narrow-line LRDs have relatively low [O iii] EWs. These different emission-line properties, along with their characteristic V-shape SEDs, suggest that narrow-line LRDs represent a special population.
@@
We also check the locations of the narrow-line LRDs and other LRDs in the Baldwin, Phillips, & Telervich (BPT) diagnostic diagrams (J. A. Baldwin et al., 1981; S. Veilleux & D. E. Osterbrock, 1987), using the measured narrow component properties. For all LRDs, the [N ii] doublets are very faint and undetected, allowing only upper limits for the [N ii]/H ratios (a 3 upper limit is adopted). For UID-16321, its H line is not detected in both the prism and grating spectrum. We put an upper limit on its [O iii]/H ratio based on the measured 3 flux upper limit of the H line in the grating spectrum. The resulting distribution of the five narrow-line LRDs on the BPT diagram is shown in Figure 7, in which the SDSS low-redshift galaxies and JWST high-redshift star-forming galaxies are plotted for comparison. The solid and dashed lines indicate the demarcation between local AGN and star-forming galaxies from L. J. Kewley et al. (2001) and G. Kauffmann et al. (2003), respectively. The narrow-line LRDs, as well as other LRDs, lie near the demarcation. Additionally, the traditional BPT diagram is likely not applicable to high-redshift objects, as high-redshift AGNs are offset from the local AGN branch and mix with the star-forming galaxy branch (e.g., D. D. Kocevski et al., 2023; Y. Harikane et al., 2023; I. Juodžbalis et al., 2025). Therefore, we are not able to distinguish AGNs from star-forming galaxies based on this diagram.
3.5 Composite Spectra
@@
4 Interpretation of Narrow-line LRDs
@@
In this section, we consider two possible interpretations of the narrow-line LRDs: they may be either compact star-forming galaxies or AGNs with low black hole masses. There is another possibility that narrow-line LRDs represent the “Type 2” counterparts of the broad-line LRDs, with the difference arising from obscuration by a dust torus. However, given the similar SEDs of the broad-line and narrow-line LRDs, this scenario appears puzzling and is not considered further in this section. We will briefly revisit it in Section 5.3.
4.1 Purely Star-forming Galaxies
@@
The presence of broad Balmer lines in LRDs is the key evidence of AGN activity. Conversely, the absence of a broad H component suggests that the AGN contribution in these narrow-line LRDs is likely negligible. The extended UV emission detected in some of the narrow-line LRDs (see Section 3.3) also supports significant galaxy contribution. Therefore, narrow-line LRDs can be purely star-forming galaxies with a special V-shape continuum.
4.1.1 Stellar Masses and Star Formation Rates
@@
To estimate their physical properties, including stellar mass and star formation rate (SFR), we perform SED modeling for the narrow-line LRDs using CIGALE and (v2025, D. Burgarella et al., 2005; M. Boquien et al., 2019) and Bagpipe (A. C. Carnall et al., 2018, 2019), assuming that the whole SED is produced by galaxy. We employ two different codes with distinct configurations to explore the redundancy of the best-fit SED solutions. Each source’s redshift is fixed at the spectroscopic redshift provided by DJA.
@@
For CIGALE, we assume a delayed- star formation history (SFH), where and . We also allow an optional late starburst in the latest 20. We use the G. Bruzual & S. Charlot (2003) stellar population synthesis models and adopt the G. Chabrier (2003) initial mass function (IMF). Nebular continuum and emission lines are included, allowing a metallicity range of [0.2, 1] and an ionization parameter . We adopt the SMC extinction curve (Y. C. Pei, 1992) for nebular emission and the modified D. Calzetti et al. (2000) attenuation curve for the stellar continuum, with the power-law slope . We add a relative error of 5% in quadrature to the uncertainties of the fluxes.
@@
For Bagpipe, we use the synthetic templates from G. Bruzual & S. Charlot (2003) with a Kroupa IMF, a stellar mass cut-off of , and nebular emission computed self-consistently with CLOUDY (G. J. Ferland et al., 2013) with . We assume a delayed- SFH, where and . We also allow an optional late starburst in the latest 20. The allowed metallicity range is [0.2, 1]. We adopt the double-component attenuation law of S. Charlot & S. M. Fall (2000) with the power-law slope . The low-resolution prism spectrum and the multi-band photometric data are fitted simultaneously in the Bagpipe fitting. We adopt a flexible calibration scheme to account for potential discrepancies in relative flux calibration between spectra and photometry. Specifically, we model the spectroscopic flux calibration uncertainties by fitting a third-order Chebyshev polynomial to the spectrum during the fitting process. This allows us to capture effects such as aperture mismatch and wavelength-dependent calibration issues. To deal with potentially underestimated errors, we also add a noise model component to fit a multiplicative factor to all of the spectroscopic uncertainties.
@@
Both CIGALE and Bagpipes yield reasonable best-fit SED models. As an example, the CIGALE best-fit result for UID-12329 is shown in Figure 9(a). The SED modeling result suggests that the broad-band SED of a narrow-line LRD can be explained as a galaxy with a high stellar mass and SFR. It shows strong dust attenuation, and the characteristic V-shape is caused by the Balmer break and differential attenuation at rest-frame UV and optical wavelengths. This is consistent with the initial interpretation of LRDs (e.g., I. Labbé et al., 2023; H. B. Akins et al., 2023; P. G. Pérez-González et al., 2024), which is expected since the broad-band SED properties of broad-line and narrow-line LRDs are similar.
@@
The inferred stellar masses and SFRs of the narrow-line LRDs are listed in Table 2 and plotted in Figure 10. Star-forming galaxies from K. Nakajima et al. (2023) and M. Curti et al. (2024), and broad-line LRDs are also shown for comparison. The narrow-line LRDs are located at the large stellar mass and SFR end compared to the normal star-forming galaxies. The stellar masses and SFRs inferred from the CIGALE and Bagpipe show deviations of dex, with the CIGALE results generally have larger stellar masses and SFRs. Such differences are commonly seen in galaxy-only SED modeling of LRDs (e.g., H. B. Akins et al., 2023). The discrepancy is possibly driven by different prescriptions for the SFH, dust attenuation law, IMF, and the inclusion of spectrum information in the Bagpipe fitting. Nonetheless, assuming either of the two measurement values does not alter any of the conclusions presented in this work.
@@
4.1.2 Comparison of UV-based and H-based SFRs
In this subsection, we assume that the whole SEDs of the LRDs are from galaxies and compare the SFRs derived from the UV continuum emission and line emission. The UV emission directly traces the photospheric emission of young stars, tracing recent star formation over the past 10–200 Myr. The H line emission traces ionized gas surrounding massive young stars, providing a nearly instantaneous measure of the SFR by probing stars with lifetimes of 3–10 Myr. They are two of the most commonly used tracers of SFR and are both sensitive to the recent star formation. We adopt measured in Section 3.2 and at rest-frame 1550 Å from the best-fit SED model obtained using CIGALE, as described in Section 4.1.1. For broad-line LRDs, we use the narrow-line component from the two-Gaussian fitting to calculate . The SFRs are derived using the calibrations from C.-N. Hao et al. (2011) and E. J. Murphy et al. (2011), as summarized in R. C. Kennicutt & N. J. Evans (2012):

### v2

@@ III.2 Selection of Narrow-line LRDs
A broad Gaussian component (FWHM 1000 ) and a narrow Gaussian component (FWHM 1000 ) at the same centroid velocity.
@@
The [N ii] lines are added if needed, and they are forced to have the same width as the narrow component of H. We do not tie the FWHM of [O iii], H, and H lines, since we want to check whether they show different line widths later. This choice does not impact the identification of narrow-line LRD candidates in the current step. To compare the quality of these two types of fitting, we calculate the Bayesian Information Criterion (BIC) parameter, defined as (2007MNRAS.377L..74L):
@@
where and correspond to the two-component fitting and the single-component fitting, respectively. Based on the above procedure, we identify five narrow-line LRDs with no broad H line component from the 32 sources. We also attempt to fit the H lines of these five sources using two narrow Gaussian components (FWHM 1000 ). The resultant BIC value still favors the single narrow component model in each case. The fraction of narrow-line LRDs in our sample is . This fraction may be biased by the complex selection process of the NIRSpec observations. On the other hand, we argue that the photometric selection is not sensitive to the line width, and thus the narrow-line LRD fraction is unlikely affected significantly.
@@ III.3 Morphology
The NIRCam images of the five narrow-line LRDs in Figure 4 show that they are very compact. To quantitatively measure their morphology and detect their potential extended emission, we fit their multi-band JWST images (Figure 5) using GalfitM (2013MNRAS.430..330H; 2013MNRAS.435..623V), a multi-band version of the two-dimensional image fitting code Galfit (2002AJ....124..266P; 2010AJ....139.2097P). We first cut images for each band from the background-subtracted mosaic images. The corresponding error cutout images are used as input sigma images. To ensure that the best-fit value is close to unity, we scale each sigma image by a factor of – such that its median background pixel value matches the standard deviation of the background pixels in the science image. Bands with no significant source detection are excluded during the fitting. For UID-16321, F090W and F115W are excluded; for UID-20547, F090W, F115W, and F150W are excluded; and for UID-8219 and UID-22015, F090W is excluded.
@@
It is still unclear whether LRDs are AGN-dominated or galaxy-dominated, particularly in the case of narrow-line LRDs. Therefore, we perform three types of model fitting: (1) a single PSF model, representing an AGN-dominated morphology; (2) a single Sérsic model, representing a pure galaxy; and (3) a combination of a PSF and a Sérsic component, representing an AGN and its host galaxy. The PSF models used are those constructed in Section II.1. We fit a region centered on the source. During the fitting procedure, the axis ratio, half-light radius, and Sérsic index of the galaxy component are allowed to vary following a -order Chebyshev polynomial, while the position angle is allowed to vary following a -order Chebyshev polynomial. These flexibilities serve to accommodate the morphological difference between long and short wavelengths, as these sources are often more extended in shorter wavelengths.
@@
For UID-16321, UID-20547, and UID-22015, their multi-band JWST image can be well fitted by a single PSF model. UID-16321 and UID-20547 are very faint in the short wavelengths, so only their long-wavelength data (where LRDs typically appear unresolved) are used to measure the morphology. For UID-20547, it F277W band residual of the PSF fitting suggests weak extended emission. For UID-8219 and UID-12329, their short-wavelength images cannot be well fitted by a single PSF model, revealing clear extended emission. UID-8219 can be sufficiently fitted with a Sérsic component, and an additional PSF component slightly decreases the value from 1.13 to 1.09. In case of UID-12329, a single Sérsic model also fails to provide a good fit, with significant residuals in the short-wavelength bands. We thus adopt a PSF+Sérsic model that yields a reasonable result with value of 1.24. Figure 5 shows the PSF+Sérsic fitting result for UID-12329. The fraction of the Sérsic component decreases toward longer wavelengths. The Sérsic component of UID-12329 is slightly offset from the central point sources; such offset extended emission are also seen in other LRDs (e.g., 2025ApJ...983...60C; 2025arXiv250503183C).
III.4 Emission-line Properties
@@
Using the emission line quantities measured in Section III.2, we analyze the emission line properties of the narrow-line LRDs and compare them with those of normal star-forming galaxies. We adopt galaxy samples from 2023ApJS..269...33N and 2024A&A...684A..75C and measure their properties using the same method as applied to our sources. Figure 6 presents a comparison of the intrinsic FWHM, luminosity, and equivalent width (EW) of the [O iii], H, and H lines among narrow-line LRDs, broad-line LRDs, and the star-forming galaxies. For broad-line LRDs, we only use the narrow components of the H and H lines for comparison. UID-16321 does not have medium- or high-resolution spectral coverage of the [O iii] line, thus its values are not available. We measure its from the prism spectrum. The comparison of the FWHM between the [O iii] and H lines in Figure 6(c) does not account for the LSF effect, as these two lines have similar wavelengths and spectral resolutions.
@@
Figure 6 suggests that narrow-line LRDs may exhibit different emission-line properties compared to typical star-forming galaxies. NIRSpec grating spectrum typicall has higher spectral resolutions at longer wavelengths. After correcting the LSF effect, is consistent with (Figure 6a). For the narrow-line LRDs, three of them have slightly larger than . On the other hand, three LRDs have medium-resolution spectra covering H and they all show consistent and . UID-22015 shows tentative larger but its measurement error is . The observed of narrow-line LRDs also tend to be larger compared with the star-forming galaxies with similar as shown in Figure 6(b). We note the line luminosities here are not corrected for dust attenuation. The larger H luminosities of narrow-line LRDs may simply reflect higher dust extinction, as [O iii] is more strongly attenuated than H. However, only mild dust extinction is expected for LRDs (e.g., 2025arXiv250518873C; 2025arXiv250522600C; 2025arXiv250302059S). Conversely, the narrow-line LRDs generally exhibit smaller H EWs and larger [O iii] EWs compared with star-forming galaxies, although two narrow-line LRDs have relatively low [O iii] EWs. These different emission-line properties, along with their characteristic V-shape SEDs, suggest that narrow-line LRDs represent a special population.
@@
We also check the locations of the narrow-line LRDs and other LRDs in the Baldwin, Phillips, & Telervich (BPT) diagnostic diagrams (1981PASP...93....5B; 1987ApJS...63..295V), using the measured narrow component properties. For all LRDs, the [N ii] doublets are very faint and undetected, allowing only upper limits for the [N ii]/H ratios (a 3 upper limit is adopted). For UID-16321, its H line is not detected in both the prism and grating spectrum. We put an upper limit on its [O iii]/H ratio based on the measured 3 flux upper limit of the H line in the grating spectrum. The resulting distribution of the five narrow-line LRDs on the BPT diagram is shown in Figure 7, in which the SDSS low-redshift galaxies and JWST high-redshift star-forming galaxies are plotted for comparison. The solid and dashed lines indicate the demarcation between local AGN and star-forming galaxies from 2001ApJ...556..121K and 2003MNRAS.346.1055K, respectively. The narrow-line LRDs, as well as other LRDs, lie near the demarcation. Additionally, the traditional BPT diagram is likely not applicable to high-redshift objects, as high-redshift AGNs are offset from the local AGN branch and mix with the star-forming galaxy branch (e.g., 2023ApJ...954L...4K; 2023ApJ...959...39H; 2025arXiv250403551J). Therefore, we are not able to distinguish AGNs from star-forming galaxies based on this diagram.
III.5 Composite Spectra
@@
IV Interpretation of Narrow-line LRDs
@@
In this section, we consider two possible interpretations of the narrow-line LRDs: they may be either compact star-forming galaxies or AGNs with low black hole masses. There is another possibility that narrow-line LRDs represent the “Type 2” counterparts of the broad-line LRDs, with the difference arising from obscuration by a dust torus. However, given the similar SEDs of the broad-line and narrow-line LRDs, this scenario appears puzzling and is not considered further in this section. We will briefly revisit it in Section V.3.
IV.1 Purely Star-forming Galaxies
@@
The presence of broad Balmer lines in LRDs is the key evidence of AGN activity. Conversely, the absence of a broad H component suggests that the AGN contribution in these narrow-line LRDs is likely negligible. The extended UV emission detected in some of the narrow-line LRDs (see Section III.3) also supports significant galaxy contribution. Therefore, narrow-line LRDs can be purely star-forming galaxies with a special V-shape continuum.
IV.1.1 Stellar Masses and Star Formation Rates
@@
To estimate their physical properties, including stellar mass and star formation rate (SFR), we perform SED modeling for the narrow-line LRDs using CIGALE and (v2025, 2005MNRAS.360.1413B; 2019A&A...622A.103B) and Bagpipe (2018MNRAS.480.4379C; 2019MNRAS.490..417C), assuming that the whole SED is produced by galaxy. We employ two different codes with distinct configurations to explore the redundancy of the best-fit SED solutions. Each source’s redshift is fixed at the spectroscopic redshift provided by DJA.
@@
For CIGALE, we assume a delayed- star formation history (SFH), where and . We also allow an optional late starburst in the latest 20. We use the 2003MNRAS.344.1000B stellar population synthesis models and adopt the 2003PASP..115..763C initial mass function (IMF). Nebular continuum and emission lines are included, allowing a metallicity range of [0.2, 1] and an ionization parameter . We adopt the SMC extinction curve (1992ApJ...395..130P) for nebular emission and the modified 2000ApJ...533..682C attenuation curve for the stellar continuum, with the power-law slope . We add a relative error of 5% in quadrature to the uncertainties of the fluxes.
@@
For Bagpipe, we use the synthetic templates from 2003MNRAS.344.1000B with a Kroupa IMF, a stellar mass cut-off of , and nebular emission computed self-consistently with CLOUDY (2013RMxAA..49..137F) with . We assume a delayed- SFH, where and . We also allow an optional late starburst in the latest 20. The allowed metallicity range is [0.2, 1]. We adopt the double-component attenuation law of 2000ApJ...539..718C with the power-law slope . The low-resolution prism spectrum and the multi-band photometric data are fitted simultaneously in the Bagpipe fitting. We adopt a flexible calibration scheme to account for potential discrepancies in relative flux calibration between spectra and photometry. Specifically, we model the spectroscopic flux calibration uncertainties by fitting a third-order Chebyshev polynomial to the spectrum during the fitting process. This allows us to capture effects such as aperture mismatch and wavelength-dependent calibration issues. To deal with potentially underestimated errors, we also add a noise model component to fit a multiplicative factor to all of the spectroscopic uncertainties.
@@
Both CIGALE and Bagpipes yield reasonable best-fit SED models. As an example, the CIGALE best-fit result for UID-12329 is shown in Figure 9(a). The SED modeling result suggests that the broad-band SED of a narrow-line LRD can be explained as a galaxy with a high stellar mass and SFR. It shows strong dust attenuation, and the characteristic V-shape is caused by the Balmer break and differential attenuation at rest-frame UV and optical wavelengths. This is consistent with the initial interpretation of LRDs (e.g., 2023Natur.616..266L; 2023ApJ...956...61A; 2024ApJ...968....4P), which is expected since the broad-band SED properties of broad-line and narrow-line LRDs are similar.
@@
The inferred stellar masses and SFRs of the narrow-line LRDs are listed in Table 2 and plotted in Figure 10. Star-forming galaxies from 2023ApJS..269...33N and 2024A&A...684A..75C, and broad-line LRDs are also shown for comparison. The narrow-line LRDs are located at the large stellar mass and SFR end compared to the normal star-forming galaxies. The stellar masses and SFRs inferred from the CIGALE and Bagpipe show deviations of dex, with the CIGALE results generally have larger stellar masses and SFRs. Such differences are commonly seen in galaxy-only SED modeling of LRDs (e.g., 2023ApJ...956...61A). The discrepancy is possibly driven by different prescriptions for the SFH, dust attenuation law, IMF, and the inclusion of spectrum information in the Bagpipe fitting. Nonetheless, assuming either of the two measurement values does not alter any of the conclusions presented in this work.
@@
IV.1.2 Comparison of UV-based and H-based SFRs
In this subsection, we assume that the whole SEDs of the LRDs are from galaxies and compare the SFRs derived from the UV continuum emission and line emission. The UV emission directly traces the photospheric emission of young stars, tracing recent star formation over the past 10–200 Myr. The H line emission traces ionized gas surrounding massive young stars, providing a nearly instantaneous measure of the SFR by probing stars with lifetimes of 3–10 Myr. They are two of the most commonly used tracers of SFR and are both sensitive to the recent star formation. We adopt measured in Section III.2 and at rest-frame 1550 Å from the best-fit SED model obtained using CIGALE, as described in Section IV.1.1. For broad-line LRDs, we use the narrow-line component from the two-Gaussian fitting to calculate . The SFRs are derived using the calibrations from 2011ApJ...741..124H and 2011ApJ...737...67M, as summarized in 2012ARA&A..50..531K: